# Open Vocabulary Monocular 3D Object Detection

### v1

@@ Abstract
In this work, we pioneer the study of open-vocabulary monocular 3D object detection, a novel task that aims to detect and localize objects in 3D space from a single RGB image without limiting detection to a predefined set of categories. We formalize this problem, establish baseline methods, and introduce a class-agnostic approach that leverages open-vocabulary 2D detectors and lifts 2D bounding boxes into 3D space. Our approach decouples the recognition and localization of objects in 2D from the task of estimating 3D bounding boxes, enabling generalization across unseen categories. Additionally, we propose a target-aware evaluation protocol to address inconsistencies in existing datasets, improving the reliability of model performance assessment. Extensive experiments on the Omni3D dataset demonstrate the effectiveness of the proposed method in zero-shot 3D detection for novel object categories, validating its robust generalization capabilities. Our method and evaluation protocols contribute towards the development of open-vocabulary object detection models that can effectively operate in real-world, category-diverse environments.
@@ 1 Introduction
Recognizing objects from a single image has been a longstanding task in computer vision, with broad applications in robotics and AR/VR. Over recent decades, 2D object detection — identifying and localizing objects within a 2D image plane — has achieved substantial progress, driven by advancements in deep learning techniques [16, 44, 20, 9] and large annotated datasets [31, 19, 49]. However, recognizing only a fixed set of objects is limiting, given the vast diversity of objects in real-world settings; detecting objects solely in 2D space is also insufficient for most practical tasks, as the world and its objects exist in 3D space.
@@
To address these challenges, one line of recent research has focused on open-vocabulary 2D object detection [32, 12, 55, 65] (Fig. 1b) to identify objects beyond a fixed set of categories. Another line explores the monocular 3D detection task [5, 47, 53, 37] (Fig. 1c), which extends detection capabilities from 2D to 3D space. Despite the vast research in these two lines of research, the intersection of these two areas — open-vocabulary monocular 3D detection, referred to as OVMono3D (Fig. 1d) — remains largely unexplored.
In this work, we aim to fill this gap. The OVMono3D task involves detecting and localizing objects of any categories in the metric 3D space, including novel categories unseen during training. We identify two primary challenges in this task: first, the lack of a well-defined task setting and standardized evaluation protocol, which complicates consistent model assessment; and second, the scarcity and limited quality of labeled 3D data, as annotating detailed 3D bounding boxes is labor-intensive (Fig. 2 and Fig. 3).
To address these challenges, we first propose a formal definition of the task and a simple protocol to mitigate the impact of annotation issues (e.g., missing labels) in evaluations (Sec. 3). Our evaluation metric enables direct performance assessment without exhaustive 3D annotations, which is especially useful for efficient real-world evaluation of a detection system, where annotating every object is impractical and error-prone.
One unique challenge in OVMono3D is the lack of large-scale 3D datasets with high-quality annotations (Fig. 2). For this reason, we adopt an approach that decouples the tasks of recognizing and localizing objects in 2D from estimating their 3D bounding boxes. We first consider a simple training-free method which un-projects 2D detections from off-the-shelf open-vocabulary 2D detectors [12, 32] into 3D via geometric principles, referred to as OVMono3D-GEO (Fig. 4a). Despite of its simplicity and effectiveness in general cases, we find this approach lacks of robustness to object occlusion.
To tackle this issue, we then introduce a data-driven approach, OVMono3D-LIFT, which learns to lift 2D detections to 3D (Fig. 4b). We comprehensively explore its design space, examining choices such as network backbones, 2D base detectors, and dataset size. Our experiments show that OVMono3D-LIFT achieves state-of-the-art performance in detecting novel objects in 3D, and demonstrates strong generalization on in-the-wild images (Fig. 7).
We summarize our contributions as follows:
We conduct the first comprehensive study on open-vocabulary monocular 3D object detection, formally define the task and evaluation metrics, and extensively explore the design space of several baseline methods.
The proposed OVMono3D-LIFT effectively decouples the tasks of 2D recognition and localization from 3D spatial estimation, achieving superior performance compared to alternative baseline methods.
We provide an in-depth analysis of design choices and demonstrate the effectiveness of our approach in zero-shot monocular 3D detection on in-the-wild images.
@@ 2 Related Work
The goal of this task is to recognize and localize objects in 2D images beyond a fixed set of predefined categories. Leveraging large-scale 2D datasets [49, 25, 39], this field has seen considerable advancement. Some approaches [57, 18, 36, 35] employ pre-trained vision-language models [24, 41], using frozen text features to detect novel categories. Other methods [28, 60, 32, 12] are pre-trained on extensive detection, grounding, and caption data to align region-text features. For example, Grounding DINO [32] incorporates grounded pre-training with cross-modal feature fusion, while YOLO-World [12] uses region-text contrastive loss and re-parameterization to enhance both accuracy and efficiency. However, such large-scale annotations are costly in 3D detection; thus, we investigate methods to adapt existing open-vocabulary 2D detectors for 3D detection.
@@
This task seeks to identify objects from any category in 3D, including those unseen during training. Prior research [33, 34, 7, 8, 58, 70, 63, 61, 40] primarily focuses on 3D detection tasks with 3D point clouds as input. [33] first proposes an open-vocabulary 3D detector using image-level class supervision from ImageNet1K [48]. [34] utilizes 2D bounding boxes from a pre-trained 2D detector [68] to build pseudo-3D label for object localization. [58] leverages various 2D foundation models to enhance the performance of 3D open-vocabulary detection. More recently, OV-Uni3DETR [54] proposes a multi-modal open-vocabulary 3D detector that accommodates both point clouds and images as the input. In contrast, our work focuses on the monocular 3D detection task that only requires RGB images as input and does not assume the availability of point cloud data at training or inference phases.
@@
refers to the task of identifying and localizing objects within a scene using 3D bounding boxes derived from single-view images. Early research in this domain primarily targeted either outdoor [11, 66, 67, 52, 10, 64, 22, 53] or indoor [13, 23, 37, 46] environments, focusing on specialized applications such as autonomous driving in urban areas and room layout estimation. The release of the extensive Omni3D dataset enabled Cube R-CNN [5] to pioneer unified monocular 3D object detection, allowing detectors to be effectively applied across multiple scene types and significantly enhancing their generalization capabilities. Following this, UniMODE [30] introduced the first successful BEV-based monocular 3D object detector capable of operating in both indoor and outdoor settings, demonstrating increased versatility in diverse environments. However, most of these approaches suffer from limited generalizability, and even advanced models like Cube R-CNN [5] and UniMODE [30] are constrained by a closed vocabulary, restricting their ability to recognize or detect object classes that were not included during training. To address this limitation, our work focuses on exploring the potential of monocular open-vocabulary 3D detection.
3 Open Vocabulary Monocular 3D Detection
3.1 Task Definition
Formally, given an input image and a text prompt consisting of all possible category names from an open vocabulary , the objective of this task is to predict a set of objects , where:
is the predicted category label for object .
is the parameters of the 3D bounding box in the camera coordinate system, defined in metric space.
The vocabulary is divided into two subsets: base categories , which are seen during training, and novel categories , which are unseen during training and only evaluated during inference. The model is trained to predict the 3D bounding boxes for objects in and generalize to without additional fine-tuning.
3.2 Evaluation Metrics
@@
Mean Average Precision (AP) based on Intersection over Union (IoU) is the widely used standard metric for closed-vocabulary 3D object detection tasks [5, 30]. However, directly applying traditional evaluation protocols to OVMono3D presents challenges due to common issues in 3D dataset annotations:
Naming ambiguity: Objects may be labeled with different conventions or policies during annotation (e.g., table vs. desk). Standard open-vocabulary 2D detection methods typically prompt 2D detectors with exhaustive lists of possible categories, which can lead to correct predictions with class names that do not align with the dataset annotations, especially for vaguely defined or overlapping categories, as shown in Fig. 3b.
@@
To address these issues, we propose a customized evaluation approach that considers only the categories with ground-truth labeled instances in each image. Specifically, instead of providing the 2D detector with an exhaustive list of possible categories, we only prompt it with category names that exist in the annotations for each image. Since human annotators typically label all instances of the same category within a single image, categories with ground-truth annotations are likely to be fully annotated.
Our target-aware evaluation prevents the negative effects of missing annotations on metrics by focusing only on categories present in the ground-truth annotations. It also excludes irrelevant categories that could distract the 2D detector. As shown in Fig. 3c, by concentrating on relevant categories, this approach mitigates the impact of missing annotations and reduces category ambiguity, enabling a more accurate assessment of the model’s 3D detection capabilities.
4 Methodology
Our approach builds on two state-of-the-art frameworks: Cube R-CNN [5] for monocular 3D detection and Grounding DINO [32] for open-vocabulary (OV) 2D detection. In this section, we first provide a brief overview of these two frameworks (Sec. 4.1); we then introduce our two proposed methods for addressing OVMono3D (Sec. 4.2 - 4.3).
4.1 Preliminaries
Cube R-CNN [5] is a state-of-the-art monocular 3D detection model, which is trained from a large-scale 3D dataset (i.e., Omni3D). It extends Faster R-CNN [44] with a 3D cube head. Using 2D region proposals as input, the cube head employs ROI poolers to extract local image features and then predicts 3D bounding boxes with MLPs. Each 3D bounding box is defined by its geometric parameters:
Position: , the 3D coordinates of the object’s center in metric space.
Dimensions: , representing width, height, and length in meters.
Orientation: Represented by a continuous 6D allocentric rotation [69].
The training objective of Cube R-CNN is defined as:
where includes the classification and bounding box regression losses from the 2D detection head [44], is the loss from the 3D cube head, and denotes the uncertainty score. The 3D loss consists of disentangled losses for each 3D attribute [50]:
where represents variable groups for the 2D center shift, depth, dimensions, and rotation. Each disentangled loss isolates the error of a specific group by substituting other predicted variables with ground-truth values when constructing the predicted 3D bounding box . The holistic loss compares the predicted 3D bounding box with the ground truth using the Chamfer Loss:
Grounding DINO [32] is a leading framework for OV 2D detection, which combines the Transformer-based DINO detector [59] with grounded pretraining. We adopt a pretrained Grounding DINO as our default OV 2D detection model due to its superior performance and strong zero-shot generalization capabilities. An ablation study on the effect of different base 2D detection models on our approach is provided in Sec. 5.
4.2 OVMono3D-GEO: Geometric 2D Unprojection
To perform OV 3D detection from a single image, a simple approach is to unproject 2D detections into 3D using geometric principles. Specifically, given an input image , a text prompt , and 2D bounding boxes with category labels predicted by an OV 2D detector [32], the method proceeds as follows.
For each detected object , an instance segmentation mask is obtained using a segmentation model (e.g., SAM [26]), and a metric depth estimation model (e.g., Depth Pro [4]) generates the depth map . Pixels within are unprojected into 3D space using the camera intrinsic matrix to form a point cloud . Here, and denote the pixel coordinates in the image, and each 3D point is computed as:
To estimate the 3D bounding box parameters , Principal Component Analysis (PCA) determines the orientation , dimensions , and centroid of . DBSCAN [14] removes outliers, improving robustness against noisy masks and depth predictions.
This geometric approach reconstructs 3D bounding boxes from 2D detections without relying on learned 3D attributes, providing a non-learning-based benchmark for evaluating OV 3D object detection models. Its accuracy depends on the quality of and , and is sensitive to occlusions, which may cause incomplete point clouds and inaccuracies in the reconstructed bounding boxes.
4.3 OVMono3D-LIFT: Class-Agnostic 2D Lifting
Given the limitations of geometric-based methods, we propose leveraging monocular 3D annotations to build data-driven models. As depicted in Fig. 2, 3D annotations are far fewer than 2D annotations. To address this, we utilize off-the-shelf 2D data-driven priors to enhance 3D performance. Specifically, we introduce OVMono3D-LIFT, which decouples OV monocular 3D detection into two stages: (1) recognizing and localizing objects in 2D using off-the-shelf OV 2D detectors, and (2) class-agnostically lifting the 2D bounding boxes to 3D cuboids.
Given an image , a text prompt , and 2D bounding boxes with category labels predicted by an OV 2D detector [32], OVMono3D-LIFT employs a pretrained Vision Transformer (ViT-B/14) backbone to extract an image feature map , where is the feature dimension. A Simple Feature Pyramid (SFP) [29] module generates three feature maps with scaling factors 0.5, 1, and 2, forming a multi-scale hierarchical pyramid. The predicted 2D bounding boxes are used as regions of interest for ROI pooling, extracting local object features of size . These features are fed into a cube head, as described in Sec. 4.1, to predict 3D attributes. Unlike [5], our attribute prediction head is class-agnostic, while their method employs class-specific layers and per-class average size statistics, which limit generalization to open vocabularies. The training objective is identical to Cube R-CNN, as described in Sec. 4.1.
Baseline Methods.
An alternative data-driven approach is to directly extend OV 2D detectors to 3D. Specifically, we use the output features from the final decoder layer of Grounding DINO [32] as object representations, which are then processed by an MLP to predict 3D attributes. While straightforward, this method does not leverage pretrained image encoders (e.g. DINOv2 [38]). Therefore, we use it as a baseline for comparison.
We evaluate two variants of this method: (1) G-DINO-3D (FZ), where only the additional MLP is trainable and other parameters are frozen, and (2) G-DINO-3D (FT), where we fine-tune all network parameters. Both variants are initialized with the pretrained weights of Grounding DINO and follow the same training objective as Cube R-CNN, as outlined in Sec. 4.1.

### v2

@@ Abstract
We propose and study open-vocabulary monocular 3D detection, a novel task that aims to detect objects of any categories in metric 3D space from a single RGB image. Existing 3D object detectors either rely on costly sensors such as LiDAR or multi-view setups, or remain confined to closed vocabularies settings with limited categories, restricting their applicability. We identify two key challenges in this new setting. First, the scarcity of 3D bounding box annotations limits the ability to train generalizable models. To reduce dependence on 3D supervision, we propose a framework that effectively integrates pretrained 2D and 3D vision foundation models. Second, missing labels and semantic ambiguities (e.g., table vs. desk) in existing datasets hinder reliable evaluation. To address this, we design a novel metric that captures model performance while mitigating annotation issues. Our approach achieves state-of-the-art results in zero-shot 3D detection of novel categories as well as in-domain detection on seen classes. We hope our method provides a strong baseline and our evaluation protocol establishes a reliable benchmark for future research.
@@ 1 Introduction
Recognizing objects from a single image has been a longstanding task in computer vision, with broad applications in robotics and AR/VR. Over recent decades, 2D object detection — identifying and localizing objects within a 2D image plane — has achieved substantial progress, driven by advancements in deep learning techniques [girshick2015fast, ren2016faster, he2017mask, carion2020detr] and large annotated datasets [lin2014microsoft, gupta2019lvis, shao2019objects365]. However, recognizing only a fixed set of objects is limiting, given the vast diversity of objects in real-world settings; detecting objects solely in 2D space is also insufficient for most practical tasks, as the world and its objects exist in 3D space.
@@
To address these limitations, one line of recent research has focused on open-vocabulary 2D object detection [liu2023grounding, cheng2024yolo, wu2024clim, zhou2022extract, wu2023open] (Fig. 1b) to identify objects beyond a fixed set of categories. Another line explores the monocular 3D detection task [brazil2023omni3d, rukhovich2022imvoxelnet, wang2022probabilistic, nie2020total3dunderstandingjointlayoutobject] (Fig. 1c), which extends detection capabilities from 2D to 3D space. Despite the vast research in these two lines of research, the intersection of these two areas — open-vocabulary monocular 3D detection, referred to as OVMono3D (Fig. 1d) — remains largely unexplored. In this work, we aim to fill this gap. The OVMono3D task involves detecting and localizing objects of any categories in the metric 3D space, including novel categories unseen during training.
One of the central challenges in OVMono3D lies in generalizable model training, as large-scale 3D datasets with high-quality annotations are scarce. Unlike open-vocabulary 2D detection, where abundant training data exists, 3D datasets are limited in scale (Fig. 2). To address this, we adopt a decoupled strategy that separates 2D recognition and localization from 3D bounding box estimation. Specifically, we explore two complementary approaches: (1) OVMono3D-GEO, a simple training-free baseline that unprojects 2D detections from off-the-shelf open-vocabulary detectors [cheng2024yolo, liu2023grounding] into 3D using geometric principles (Fig. 3a). While requiring no 3D annotations and achieving reasonable performance, this approach is sensitive to occlusion and degrades in cluttered scenes; (2) OVMono3D-LIFT, a data-driven method that integrates multiple vision foundation models (e.g., depth estimation [piccinelli2024unidepth], 2D open-vocabulary detection [cheng2024yolo, liu2023grounding], and general image encoder [oquab2023dinov2]) and learns to lift 2D detections into 3D (Fig. 3b). We systematically analyze its design space, including backbone architectures, 2D base detectors, and other critical choices (Sec. 5.3). Our results highlight the importance of 3D-aware image representations and accurate depth estimation for robust performance. Compared to the geometric baseline, OVMono3D-LIFT shows significantly greater resilience to occlusion and achieves superior results across diverse, real-world scenarios.
Another central challenge is reliable model evaluation, hindered by missing annotations and semantic ambiguity. The task requires detecting all visible objects in 2D images, yet existing 3D datasets generally provide only partial annotations due to the high cost of labeling. For example, in SUN RGB-D [song2015sun], a chair visible in the RGB image may be unannotated in 3D if it was partially occluded in the depth scan, leading to false negatives in evaluation. Similarly, inconsistencies in naming conventions across datasets, such as predicting desk when the ground truth label is table, cause correct detections to be miscounted as false positives under standard metrics. To address these issues, we propose a simple evaluation protocol that reduces the impact of missing labels and semantic variation (Sec. 4). The resulting metric enables reliable assessment without exhaustive annotations and is particularly suited for in-the-wild images, where complete labeling is infeasible and error-prone.
Finally, we conduct extensive experiments and ablations to validate both our proposed approaches and evaluation metric on standard benchmarks. Our method not only surpasses concurrent open-vocabulary 3D detectors (e.g., DetAny3D [zhang2025detect]) by a large margin (Tab. 1), but also outperforms closed-vocabulary counterparts (e.g., Cube R-CNN[brazil2023omni3d], UniMODE [li2024unified3dobjectdetection]) on in-domain categories (Tab. 2), under both standard and proposed metrics. Qualitative results further highlight strong generalization to in-the-wild images (Fig. LABEL:fig:coco). We hope these insights inspire further exploration of advanced architectures for this task. Code will be released to support future research.
@@ 2 Related Work
The goal of this task is to recognize and localize objects in 2D images beyond a fixed set of predefined categories. Leveraging large-scale 2D datasets [shao2019objects365, kamath2021mdetr, ordonez2011im2text], this field has seen considerable advancement. Some approaches [zareian2021open, gu2021open, minderer2022simple, ma2022open] employ pre-trained vision-language models [jia2021scaling, radford2021learning], using frozen text features to detect novel categories. Other methods [li2022grounded, zhang2022glipv2, liu2023grounding, cheng2024yolo] are pre-trained on extensive detection, grounding, and caption data to align region-text features. For example, Grounding DINO [liu2023grounding] incorporates grounded pre-training with cross-modal feature fusion, while YOLO-World [cheng2024yolo] uses region-text contrastive loss and re-parameterization to enhance both accuracy and efficiency. However, such large-scale annotations are costly in 3D detection; thus, we investigate methods to adapt existing open-vocabulary 2D detectors for 3D detection.
@@
This task seeks to identify objects from any category in 3D, including those unseen during training. Prior research [lu2022open, lu2023open, cao2024coda, cao2024collaborative, zhang2024fm, zhu2023object2scene, zhang2022pointclip, zhang2025opensight, peng2025global, xia2024openad, peng2025glrd, Djamahl2024find, yao2025labelany3d] primarily focuses on 3D detection tasks with 3D point clouds as input. [lu2022open] first proposes an open-vocabulary 3D detector using image-level class supervision from ImageNet1K [russakovsky2015imagenet]. [lu2023open] utilizes 2D bounding boxes from a pre-trained 2D detector [zhou2022detecting] to build pseudo-3D label. [zhang2024fm] leverages various 2D foundation models to enhance the performance of 3D open-vocabulary detection. OV-Uni3DETR [wang2024ov] proposes a multi-modal open-vocabulary 3D detector that accommodates both point clouds and images as the input.
In contrast, our work focuses on the monocular 3D detection task that only requires RGB images as input and does not assume the availability of point cloud data at training or inference phase.
@@
refers to the task of identifying and localizing objects within a scene using 3D bounding boxes derived from single-view images. Early research in this domain primarily targeted either outdoor [Chen_2016_CVPR_Mono_Auto_Driving, zhou2021iafainstanceawarefeatureaggregation, zhou2019objectspoints, wang2021fcos3dfullyconvolutionalonestage, chen2021monorunmonocular3dobject, zhang2023monodetr, huang2022monodtr, wang2022probabilistic] or indoor [Dasgupta_2016_CVPR, huang2019cooperativeholisticsceneunderstanding, nie2020total3dunderstandingjointlayoutobject, rukhovich2021imvoxelnetimagevoxelsprojection, lazarow2024cubify] environments, focusing on specialized applications such as autonomous driving and room layout estimation. The extensive Omni3D [brazil2023omni3d] dataset enabled Cube R-CNN [brazil2023omni3d] and UniMODE [li2024unified3dobjectdetection] to perform unified monocular 3D object detection across multiple scene types. Recent works also explore pseudo annotation generation to expand detector vocabularies. OVM3D-Det [huang2024training] generates pseudo labels enabling detection of novel categories without human annotations. V-MIND [jhang2025v] converts large-scale 2D datasets into pseudo 3D training data to expand detection vocabulary.
However, most of these approaches suffer from limited generalizability, and even advanced models [brazil2023omni3d, li2024unified3dobjectdetection] are constrained by a closed vocabulary, restricting their ability to recognize or detect object classes that were not included during training. To address this limitation, our work focuses on exploring the potential of monocular open-vocabulary 3D detection.
Concurrent to our work, DetAny3D [zhang2025detect] develops a promptable 3D detection model using large-scale data and foundation models. This work addresses similar problems to ours through different methodologies. We compare with DetAny3D in our experiments quantitatively with extensive experiments.
3 Methodology
Our approach builds on two frameworks: Cube R-CNN [brazil2023omni3d] for closed-vocabulary monocular 3D detection and Grounding DINO [liu2023grounding] for open-vocabulary (OV) 2D detection. In this section, we first provide an overview of these frameworks (Sec. 3.1); we then introduce our proposed methods for addressing OVMono3D (Sec. 3.2 - 3.3, Fig. 3).
3.1 Preliminaries
Cube R-CNN [brazil2023omni3d] is a state-of-the-art monocular 3D detection model trained on a large-scale 3D dataset (i.e., Omni3D). It extends Faster R-CNN [ren2016faster] with a 3D cube head. Using 2D region proposals as input, the cube head employs ROI poolers to extract local image features and then predicts 3D bounding boxes with MLPs. The training objective of Cube R-CNN is defined as:
where includes the classification and bounding box regression losses from the 2D detection head [ren2016faster], is the loss from the 3D cube head, and denotes the uncertainty score. The 3D loss consists of disentangled losses for each 3D attribute [simonelli2019disentangling]:
where represents variable groups for the 2D center shift, depth, 3D dimensions (i.e., width, height, and length in meters), and 3D poses. isolates the error of a specific group by substituting other predicted variables with ground-truth values when constructing the predicted 3D bounding box , while compares the vertices of predicted 3D bounding box with the ground truth using Chamfer distance.
Grounding DINO [liu2023grounding] is a leading framework for OV 2D detection, which combines the Transformer-based DINO detector [zhang2022dino] with grounded pretraining. We adopt a pretrained Grounding DINO as our default OV 2D detection model due to its superior performance and strong zero-shot generalization capabilities. An ablation study of different base 2D detectors on our approach is provided in Sec. 5.
3.2 Training-free OVMono3D-GEO
We first establish a training-free baseline that unprojects 2D detections into 3D using geometric principles (Fig. 3a). Given an input image , a text prompt , and 2D bounding boxes with category labels predicted by an OV 2D detector [liu2023grounding], the method proceeds as follows.
3D Geometric Prediction. For each detected object , an instance segmentation mask is obtained using a segmentation model (e.g., SAM [kirillov2023segment]), and a metric depth estimation model (e.g., UniDepthv2 [piccinelli2025unidepthv2]) generates the depth map . Pixels within are unprojected into 3D space using the camera intrinsic matrix to form a point cloud . Here, and denote the pixel coordinates in the image, and each 3D point is computed as:
3D Bounding Box Generation. To estimate the 3D bounding box parameters from the point cloud , we first apply Principal Component Analysis (PCA) on to estimate the object orientation from the principal component direction. PCA determines the object’s main orientation by identifying the directions of maximum variance, enabling tight bounding box fitting. DBSCAN clustering [ester1996density] is then applied to remove outliers from noisy depth predictions and imperfect segmentation masks, ensuring robust parameter estimation. Finally, the centroid and dimensions are computed from the cleaned point cloud along the estimated orientation .
While this method does not require 3D supervision, its accuracy relies heavily on the quality of depth estimation and segmentation, and it is particularly sensitive to occlusions, where incomplete point clouds often lead to inaccurate bounding box reconstruction (see Fig. 10 in the supplementary). This approach also shares similarities with the 3D bounding box generation pipeline in OVM3D-Det [huang2024training]. However, their work employs the generated 3D boxes as pseudo labels to expand the vocabulary of training data, without directly evaluating the accuracy of the 3D box generation itself as a OVMono3D method.
3.3 Data-driven OVMono3D-LIFT
| Methods | Novel Categories in Omni3D | Novel Dataset | |||||
| AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP ↑ | |
| OVM3D-Det [huang2024training] | 6.74 / 18.30 | 11.35 / 29.69 | 5.95 / 15.87 | 0.73 / 2.31 | 9.70 / 22.99 | 5.05 / 15.63 | 0.04 / 0.05 |
| DetAny3D [zhang2025detect] | 9.22 / 21.44 | - | - | - | - | - | 11.05 / 15.71 |
| OVMono3D-GEO | 8.48 / 20.63 | 14.07 / 35.26 | 8.22 / 18.26 | 0.49 / 1.01 | 9.71 / 21.20 | 7.78 / 20.30 | 0.73 / 0.90 |
| OVMono3D-LIFT* | 7.31 / 18.61 | 11.47 / 27.95 | 7.06 / 17.62 | 0.93 / 3.48 | 11.39 / 26.14 | 4.97 / 14.31 | 6.57 / 10.21 |
| OVMono3D-LIFT | 9.61 / 24.08 | 15.00 / 37.45 | 9.39 / 23.24 | 1.41 / 4.00 | 14.71 / 30.13 | 6.7 / 20.62 | 12.88 / 16.76 |
| Method | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ |
| Closed-vocabulary methods | ||||||||||||
| SMOKE [liu2020smoke] | - | - | - | - | - | - | - | - | - | - | - | 9.60 |
| FCOS3D [wang2021fcos3dfullyconvolutionalonestage] | - | - | - | - | - | - | - | - | - | - | - | 9.80 |
| PGD [wang2022probabilistic] | - | - | - | - | - | - | - | - | - | - | - | 11.20 |
| ImVoxelNet [rukhovich2022imvoxelnet] | - | - | - | - | - | - | - | - | - | - | - | 9.40 |
| Cube R-CNN [brazil2023omni3d] | 32.47 | 29.98 | 15.23 | 7.39 | 41.72 | 53.85 | 25.51 | 9.59 | 28.32 | 12.05 | 8.50 | 23.68 |
| UniMODE [li2024unified3dobjectdetection] | 29.20 | 36.00 | 23.00 | 8.10 | 48.00 | 66.10 | 30.20 | 10.60 | 31.10 | 14.90 | 8.70 | 28.20 |
| Open-vocabulary methods | ||||||||||||
| OVM3D-Det [huang2024training] | 7.59 | 14.25 | 11.82 | 5.34 | 25.23 | 1.93 | 9.33 | 1.36 | 10.51 | 4.82 | 3.71 | 9.49 |
| DetAny3D [zhang2025detect] | 31.61 | 30.97 | 18.96 | 7.17 | 46.13 | 54.42 | - | - | - | - | - | 24.92 |
| OVMono3D-GEO | 1.22 | 3.23 | 9.99 | 3.98 | 21.03 | 9.22 | 5.98 | 0.23 | 9.94 | 1.59 | 0.15 | 6.64 |
| OVMono3D-LIFT* | 25.58 | 30.56 | 19.52 | 9.92 | 49.34 | 62.72 | 29.04 | 11.91 | 33.96 | 12.20 | 7.21 | 27.09 |
| OVMono3D-LIFT | 31.44 | 32.47 | 23.24 | 11.89 | 54.21 | 63.48 | 32.23 | 12.90 | 37.51 | 13.35 | 8.96 | 29.63 |
To overcome the limitations of the geometric-based approach, we propose OVMono3D-LIFT (Fig. 3b), which learns to estimate 3D bounding boxes from large-scale 3D annotations. This framework decouples OVMono3D into two stages: (1) recognizing and localizing objects in 2D with off-the-shelf open-vocabulary detectors, and (2) class-agnostically lifting the 2D bounding boxes into 3D cuboids. We detail the key components below.
Image Encoder. OVMono3D-LIFT employs a pretrained DINOv2 [oquab2023dinov2] to extract image features, and a Feature Pyramid module [li2022exploring] to generates multi-scale hierarchical feature pyramid. The predicted 2D bounding boxes are used as regions of interest for ROI pooling, extracting local object features of size . These features are fed into a cube head, as described in Sec. 3.1, to predict 3D attributes.
Depth Estimator. Similar to OVMono3D-GEO, the method first obtain the predicted depth map , then unproject it to a point map using the camera intrinsic matrix . ROI pooling is then performed on the point map using the 2D bounding boxes to extract local geometric information of size . This geometric information is concatenated with the visual ROI features to form geometry-informed features, which are input into the 3D cube head to predict 3D box attributes. During both training and inference, the method uses pseudo depth maps obtained from a pretrained metric depth estimator—no ground truth depth maps are required. By fusing geometric information into the model, the network can better infer object depth and dimensions, producing more accurate predictions.
Class-agnostic 3D Lifting head. Unlike Cube R-CNN [brazil2023omni3d], our 3D cube prediction head is class-agnostic, while their method employs class-specific layers and per-class average size statistics, which limit generalization to novel categories. The model trains only on base categories with the training objective identical to Cube R-CNN, as described in Sec. 3.1. Since the method also train a 2D head on base categories, the model preserves detection ability on these categories. During inference, for novel categories, an OV 2D detector (e.g., Grounding DINO [liu2023grounding]) is used to obtain 2D boxes, which serve as input to the ROI pooling module.
OVMono3D-LIFT*.
We include a model variant as a baseline. It removes the depth estimation module from the full OVMono3D-LIFT framework, relying solely on visual features from the image encoder and Feature Pyramid Network for 3D bounding box prediction. This variant serves to evaluate the contribution of geometric depth information by comparing performance against the full model that incorporates both visual and geometric features.
4 Evaluation Metrics
@@
Mean Average Precision (AP) based on Intersection over Union (IoU) is the widely used standard metric for closed-vocabulary 3D object detection tasks [brazil2023omni3d, li2024unified3dobjectdetection]. However, directly applying traditional evaluation protocols to OVMono3D presents challenges due to common issues in 3D dataset annotations:
Missing annotations: Comprehensive 3D annotation is often impractical due to high costs. Fig. 4a illustrates an example of this issue, where the book is not labeled.
Naming ambiguity: Objects may be labeled with different naming conventions or annotation policies during annotation (e.g., table vs. desk). Standard open-vocabulary 2D detection methods typically prompt 2D detectors with exhaustive lists of possible categories, which can lead to correct predictions with class names that do not align with the dataset annotations, especially for vaguely defined or overlapping categories, as shown in Fig. 4 (chair vs. sofa, vase vs. potted plant).
@@
To address these issues, we propose a simple evaluation approach that considers only the categories with ground-truth labeled instances in each image. Specifically, instead of providing the 2D detector with an exhaustive list of possible categories, we only prompt it with category names that exist in the annotations for each image. Since human annotators typically label all instances of the same category within a single image, categories with ground-truth annotations are likely to be fully annotated. Fig. 4 illustrates our approach enables more accurate model assessment despite annotation issues. For instance, the example highlights missing annotations for “books,” and naming ambiguities between “chair” and “sofa.” This validates the practicability of our evaluation under annotation issues.